OpenStudy (anonymous):

It took the construction company 63 days to complete the building. How many weeks did it take? (1 week = 7 days) A. 441 weeks B. 70 weeks C. 56 weeks D. 9 weeks

OpenStudy (anonymous):

63/7=?

OpenStudy (anonymous):

what does / mean?

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

ok so 63 divided by 7?

OpenStudy (anonymous):

If I give you the answer you won't learn.

OpenStudy (anonymous):

so its 9?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

\[63 days \times \frac{1 week}{7 days} = \frac{63days\times1week}{7days}\] Notice your "days" units cancel and you are left with weeks. \[\frac{63}{7}weeks\]

OpenStudy (anonymous):

Just a say to set this type of problem up (using dimensional analysis). You can do this for all types of unit conversions.

OpenStudy (anonymous):

That was a good way of explaining it :)

OpenStudy (anonymous):

way*

OpenStudy (anonymous):

My brain is a bit broken.

OpenStudy (anonymous):

No joke

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Its not funny

OpenStudy (anonymous):

I am sorry that your brain is broken. I am proud of you for doing math with a broken brain. That takes a lot of courage :)